Dus f Facebook × | P InSite Con ses/40133/quizzes/84164 Question 15 0/1 pts The following regression analysis is for the mile can someone help me with this problem and explain it? :)
Answer #1

15)

from above data regression equation :miles per gallon =44.678846-0.0069050641*weight

hence predicted weight=44.678846-0.0069050641*3000=23.96~ 24.0 mpg